OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 3 centimeters and a base that is an equilateral triangle with sides of 4 centimeters?

OpenStudy (anonymous):

length times area is vol, which is 36 I believe... If I have the right equation. If not, sorry :)

OpenStudy (anonymous):

Let me give you the answer options

OpenStudy (anonymous):

24cm3 12sqrt3 cm3 48cm3 8sqrt3 cm3

OpenStudy (ash2326):

Let's find area of the equilateral triangle with side 4 length of perpendicular= \(\sqrt{4^2-2^2}=\sqrt {12}=3.46 cm \) Area of equilateral triangle=\(\frac{1}{2} \times 3.46 \times 4=6.92 cm^2\) Volume = area * height= \(6.92\times 3=20.76 cm^2\)

OpenStudy (anonymous):

So do I just put my new numbers in place of the ones in the problem

OpenStudy (ash2326):

Yeah, find area of equilateral triangle and take the product of it with height

OpenStudy (anonymous):

I don't know where these values are coming from. I am not that far along in math

OpenStudy (ash2326):

Area of equilateral triangle =\(\frac{\sqrt{3} }{4}a^2\) a= side of equilateral triangle Volume= area of equilteral triangle \(\times\) height